payton:

Solve the following system of equations: x − 2y = 14 x + 3y = 9 (1, 12) (−1, −12) (12, −1) (12, 1)

justus:

x+3y=9 x-2y=14 5y=-5 y=-1 x=14+2*(-1)=12

justus:

Therefore, the answer choice appropriate would be (12 , -1).

payton:

thank you sm everyone

justus:

Of course!

justus:

You are most welcome.

payton:

i got a 100% i appreciate it

justus:

Yay!! :)

justus:

Good for you! I wish you the best.